# NURBGen: High-Fidelity Text-to-CAD Generation through LLM-Driven NURBS Modeling

### v1

@@ Introduction
Computer-Aided Design (CAD) plays a fundamental role in modern engineering, product design, and digital manufacturing workflows (cad_additive_review; cad_status_engineering). It enables precise, parametric modeling of complex mechanical and architectural components. However, creating detailed CAD models typically requires expert knowledge of professional design software-such as Onshape (https://www.onshape.com) or AutoCAD (https://www.autodesk.com/products/autocad/overview), and remains a labor-intensive and, time-consuming task.
Researchers have therefore proposed deep learning-based approaches for automatic CAD modeling from high-level inputs such as natural language (text2cad), images (img2cad), or point clouds (point2cad). Among these, text-to-CAD generation offers a simple, intuitive interface that allows designers to describe 3D objects in natural language, bypassing the need for expert modeling skills. However, nearly all prior methods (cadllama; cadgpt; text2cad) rely on design-history-based representations (deepcad; cadsignet; cadparser), where shapes are constructed via sequences of parametric operations—extrusions, and 2D sketches. While intuitive and highly editable, these methods are trained on small-scale datasets like DeepCAD (deepcad), which mostly contain low-complexity parts (e.g., cuboids, cylinders), limiting generalization in real-world scenarios.
@@
In contrast, the ABC dataset (abc), which contains over a million 3D CAD models, remains comparatively underutilized in text-to-CAD research due to two key limitations. First, ABC represents geometry in Boundary Representation (BRep) form, which lacks design history. BReps define solids using analytic surface patches—most commonly NURBS, the industry standard for their precision and parametric control. However, NURBS-based modeling is rarely explored in deep generative research due to the challenge of efficient representation (nurbot), non-differentiability of knot vectors (nurbsdiff), high parameter variability, and trimming complexity inherent to NURBS geometry. NeuroNURBS (neuronurbs) partially addresses this by learning latent codes for untrimmed NURBS surfaces via a non-autoregressive transformer VAE, but it does not support language-based generation and cannot model complex shapes due to the trimming issues. Second, ABC lacks high-quality text descriptions, making it difficult to train or evaluate text-conditioned generative models.
In this work, we present NURBGen, the first framework for generating 3D CAD models from natural language using structured, symbolic NURBS representations. Unlike prior work that learns dense latent codes (neuronurbs), we treat each NURBS surface as a language-aligned object: a sequence of tokens encoding control points, degrees, weights, and knot vectors in JSON format. This allows us to formulate text-to-CAD as a language modeling task. We fine-tune a large language model (Qwen3-4B) to map textual descriptions to these NURBS parameters, producing outputs that are editable and directly compatible to BRep format. To support this, we construct partABC, a curated dataset of more than 300k part-level CAD models from the ABC dataset, each represented as a sequence of NURBS surfaces and serialized with manageable context lengths ($\leq$ 8k tokens). We also generate high-quality natural language descriptions of the CAD models using an automatic annotation pipeline for the supervised fine-tuning task.
A key design choice to manage context length is our use of untrimmed NURBS surfaces similar to NeuroNURBS (neuronurbs). However, this introduces a limitation it cannot capture trimmed geometry precisely. To address this, we propose a hybrid symbolic representation that replaces NURBS with primitive analytic curves (e.g., circles, B-splines, arcs, and lines) to accurately model such faces. This maintains the structural format required for LLM fine-tuning and inference. Our experiments demonstrate that NURBGen can outperform the current state-of-the-art methods in high-fidelity text-to-CAD generation. Some examples are shown in Figure 1.
Our contributions can be summarized as follows
@@
CAD Generation: Earlier approaches to CAD generation primarily focused on low-level geometry tasks such as surface fitting (parsenet; point2cad), point cloud classification (pointnet; pointnet++), BRep segmentation (cadopsnet; brepgat; sharp2023), or BRep structure prediction (complexgen; brepdetnet), rather than generating fully parametric CAD models. A major shift came with DeepCAD (deepcad), which introduced a design-history-based representation where CAD models are expressed as sequences of 2D sketches and 3D operations (e.g., extrusions). This formulation enabled sequence-to-sequence modeling of CAD generation. Building on this, later works explored cross-modal CAD synthesis from point clouds (cadsignet; transcad; cadrecode), images (img2cad), natural language (cadllama; cad-instruct; text2cad; cadtranslator; cadmium; text-to-cad), or combinations thereof (cadrille; cad-mllm). While design-history representations are highly interpretable and editable, their reliance on proprietary CAD operation data presents a major bottleneck for large-scale public research. Public datasets like DeepCAD-170k (deepcad), Fusion360-8k (fusion360), and CADParser-50k (cadparser) are limited in size and complexity, often consisting of simple, synthetic parts (cadmium), which restricts generalization to real-world scenarios. Alternative approaches for CAD generation operate directly on BRep geometry (brepnet) or leverage SDF supervision (extrudenet; secadnet; caprinet). However, these methods don’t generalizes well. Recent work BrepGen (brepgen), for example, generates BRep topology including vertices, edges, and faces via a hierarchical latent diffusion model. In contrast, we represent BReps as sequences of structured NURBS surfaces, allowing us to frame text-to-CAD as a language generation task. This enables fine-tuning an LLM on partABC, which is more diverse and larger than those used in prior work.
@@
Nurbs Modeling: The adoption of analytic surfaces like NURBS in learning-based systems remained limited (nurb_survey; nurb_optimization). NURBSDiff (nurbsdiff) introduced differentiable NURBS fitting for geometry optimization and reconstruction. (drpg) proposed differentiable rendering of NURBS surfaces for inverse graphics tasks. The most relevant prior work is NeuroNURBS (neuronurbs), which encodes untrimmed NURBS surfaces using a non-autoregressive transformer autoencoder into latent vectors for supporting tasks like reconstruction or segmentation, but not text-conditioned generation. However, the exclusive use of untrimmed NURBS surfaces limits generalization, as not all CAD models can be accurately represented without trimming. To address this, we adopt a hybrid strategy: while untrimmed NURBS serve as our primary representation, we replace them with analytic primitives such as lines, arcs, and B-splines for faces where NURBS fitting fails.
LLM for 3D Generation: LLMs have been widely adopted across domains such as robotics (llm_robotics) and 3D scene understanding or grounding (pointllm; 3dllm). Their application to 3D generation is a relatively new but promising direction, as LLMs offer strong spatial priors and multimodal reasoning capabilities. A central challenge, however, lies in encoding 3D geometry into a sequential format compatible with language modeling. Recent work such as LLaMA-Mesh (llama_mesh) finetunes LLaMA (llama3) to generate mesh vertices and faces as plain text, demonstrating the potential of autoregressive text-based 3D synthesis. In the CAD domain, existing LLM-driven text-to-CAD methods primarily rely on design-history-based representations (cadllama; cadrecode; cad-mllm; llm_cad_survey). However, these methods are constrained by the scarcity and simplicity of the public datasets. In contrast, our method introduces a structured NURBS-based representation that enables symbolic, surface-level generation. This formulation aligns naturally with language modeling and allows us to leverage the large-scale and geometrically diverse partABC dataset for fine-tuning LLMs for the text-to-CAD task.
@@ 1. CAD Representation
A BRep solid models geometry as a collection of topologically connected faces, each defined by a bounded parametric surface. In modern CAD systems, these surfaces are most commonly represented using NURBS surfaces due to their ability to accurately model both analytic primitives (e.g., planes, cylinders, tori) and complex free-form geometry with high continuity and compactness. To reconstruct a BRep solid in a symbolic and editable form, it is essential to extract the full set of NURBS surface parameters for each face. Using pythonOCC  [^2] https://github.com/tpaviot/pythonocc-core, we propose a robust pipeline for converting BRep models into a fully parametric NURBS representation.
@@
We represent each face using either its extracted NURBS parameters or its analytic primitive definition, based on reconstruction quality. In practice, about 70% of faces are modeled using NURBS, while 30% fall back to analytic primitives. This hybrid representation stored in structured JSON offers a more expressive and compact alternative to the purely NURBS-based format used in NeuroNURBS (neuronurbs). While analytic primitives can represent simple geometry, they lack the flexibility to capture free-form surfaces. NURBS, on the other hand, provide a unified framework that can model both standard analytic shapes and complex free-form surfaces within a single patch—often replacing multiple primitives such as segmented arcs or partial cylinders. By combining both representations, our hybrid approach improves robustness and reduces parameter count for simpler shapes, resulting in shorter and more token-efficient inputs for LLM fine-tuning.
@@ 2. Annotation Pipeline
Supervised fine-tuning of our text-to-CAD model requires paired textual descriptions, but the ABC dataset lacks captions. To address this, we design an automated annotation pipeline using a vision–language model (VLM) to generate high-quality captions for CAD models at scale as show in Figure. 2 (middle).
Multi-View Rendering: Each BRep is first converted into a textureless triangular mesh and rendered from six viewpoints at a resolution of $512\times 512$ using Blender  [^3] https://docs.blender.org/api/3.6/index.html. Four of the camera views follow the orientation strategy proposed in (marvel), while the remaining two capture the top and bottom perspectives. To enhance geometric perception and visual clarity, we enable Blender’s Freestyle renderer to overlay clean silhouette and edge contours on each image.
Metadata Guidance for Caption Generation: High-quality captions for CAD models should go beyond simple naming and incorporate essential geometric features such as the number of through-holes, overall dimensions, surface area, and volume. Previous work like Text2CAD (text2cad) leverages minimal JSON-based design history to guide vision-language models (VLMs), while MARVEL (marvel) uses structured metadata for fine-grained 3D annotation. Building on these ideas, we extract geometric metadata that is often inaccessible to VLMs—specifically, length, width, height, surface area, volume, and the number of topological holes (genus). We compute overall dimensions by fitting an axis-aligned bounding box to the CAD geometry using OpenCascade’s Bnd_Box. Volume and surface area are obtained using OpenCascade’s built-in mass property computation (brepgprop.VolumeProperties and SurfaceProperties).
To estimate the number of through-holes, we first generate a watertight mesh from BRep. We then compute the Euler characteristic, $\chi=V-E+F$, where $V,E,F$ are the mesh vertices, edges, and faces. Using the Euler–Poincaré formula for closed 2-manifolds, the genus is given by $g=0.5\times(2-\chi)$, which corresponds to the number of topological through-holes (cibulkajoint). This metadata is then injected into the annotation prompt, which guides the VLM to generate captions with precise measurements.
Caption Generation: We use InternVL3-13B (internvl3), a multi-view VLM, which takes six rendered views of the CAD model along with the metadata-augmented annotation prompt as input. It then processes multi-view images of the CAD model simultaneously to generate a coherent and geometry-aware caption. Rather than focusing solely on object category names, we prioritize shape-centric descriptions that capture structural characteristics (”a bent tube..”,a flat cylindrical bolt with six holes..). The inclusion of dimensional metadata and hole counts further grounds the captions in precise geometric details, resulting in more informative and reliable annotations. In Figure 5, we provide some examples.
@@
where $l_{1}=0.35,l_{2}=0.3,l_{3}=0.25,l_{4}=0.1$ are selected using empirical experiment on 100 samples. token_count refers to the size of tokens after tokenizing the JSON using Qwen3 Tokenizer (qwen3), through_holes counts the number of holes that pass through the entire part, and bbox_diag is the length of the diagonal of the part’s axis-aligned bounding box. Based on $\texttt{w}(B)$, we categorize parts into simple ($\leq 0.12$), moderate (0.120.120.230.23), and complex (>0.23>0.23) tiers. From 3M extracted parts, we retain 10% simple, 50% moderate, and 40% complex models, forming the final partABC dataset of $\sim$300k high-quality samples as shown in Figure 6.
@@
Implementation Details: We fine-tune the Qwen3-4B model (qwen3) using AdamW (adamw) with a learning rate of $5{\times}10^{-5}$ and linear warm-up. LoRA (lora) is applied with rank $64$ and $\alpha{=}128$. Training runs for 180180k steps with batch size $1$ on $4{\times}$H200 GPUs over 3 days. The context window is $8192$ during training and 1414k during inference, with temperature $0.3$. On RTX 3090, the model achieves a generation throughput of $\sim 800$ tokens per second. Figure 2 (Right column) shows the finetuning task.
Baselines: We compare against strong open-source baselines for text-to-CAD generation. While recent models like CAD-LLaMA (cadllama) and CADFusion (text-to-cad) report promising results, their implementations are not publicly available. Moreover, to the best of our knowledge, there are no open-source models capable of generating NURBS-based CAD representations from text. Hence, we focus our comparison with open-source methods, including Text2CAD (text2cad), DeepCAD (deepcad), and GPT-4o. We use official pretrained weights for Text2CAD and retrain DeepCAD for 100 epochs following the Text2CAD protocol. GPT-4o is evaluated using 2-shot prompting with example caption–JSON pairs.
@@
Caption Quality: Since VLMs are prone to hallucinations (vlm_hallucinations), we perform an evaluation to assess the accuracy of our automatically generated captions. We randomly sample 1,000 captions and provide them along with the corresponding six rendered views and associated CAD metadata to GPT-4o. GPT-4o is tasked with verifying the correctness of each caption. We observe an estimated accuracy of $\sim 85\%$ across the sampled set. This indicates that our captioning pipeline produces high-quality and semantically meaningful descriptions.
@@
Despite its strong performance, our approach has certain limitations. Figure 7 illustrates a few representative failure cases. For instance, in response to complex prompts (e.g., “Two-story house with gabled roof…”), NURBGen struggles to capture fine-grained architectural structure. In rare cases, we also observe geometric artifacts such as self-intersections or topological inconsistencies, as seen in the second example. Additionally, NURBGen has difficulty reconstructing prompts that require text to be engraved within the geometry (third example).

### v2

@@ Introduction
Computer-Aided Design (CAD) plays a fundamental role in modern engineering, product design, and digital manufacturing workflows (Vido et al. 2024; Gao et al. 2015). It enables precise, parametric modeling of complex mechanical and architectural components. However, creating detailed CAD models typically requires expert knowledge of professional design software-such as Onshape (https://www.onshape.com) or AutoCAD (https://www.autodesk.com/products/autocad/overview), and remains a labor-intensive and, time-consuming task.
Researchers have therefore proposed deep learning-based approaches for automatic CAD modeling from high-level inputs such as natural language (Khan et al. 2024b), images (Chen et al. 2024), or point clouds (Liu et al. 2023). Among these, text-to-CAD generation offers a simple, intuitive interface that allows designers to describe 3D objects in natural language, bypassing the need for expert modeling skills. However, nearly all prior methods (Li et al. 2025; Kapsalis 2024; Khan et al. 2024b) rely on design-history-based representations (Wu et al. 2021; Khan et al. 2024a; Zhou et al. 2023), where shapes are constructed via sequences of parametric operations—extrusions, and 2D sketches. While intuitive and highly editable, these methods are trained on small-scale datasets like DeepCAD (Wu et al. 2021), which mostly contain low-complexity parts (e.g., cuboids, cylinders), limiting generalization in real-world scenarios.
@@
In contrast, the ABC dataset (Koch et al. 2019), which contains over a million 3D CAD models, remains comparatively underutilized in text-to-CAD research due to two key limitations. First, ABC represents geometry in Boundary Representation (BRep) form, which lacks design history. BReps define solids using analytic surface patches—most commonly NURBS, the industry standard for their precision and parametric control. However, NURBS-based modeling is rarely explored in deep generative research due to the challenge of efficient representation (Yang et al. 2024), non-differentiability of knot vectors (Prasad et al. 2022), high parameter variability, and trimming complexity inherent to NURBS geometry. NeuroNURBS (Fan et al. 2024) partially addresses this by learning latent codes for untrimmed NURBS surfaces via a non-autoregressive transformer VAE, but it does not support language-based generation and cannot model complex shapes due to the trimming issues. Second, ABC lacks high-quality text descriptions, making it difficult to train or evaluate text-conditioned generative models.
In this work, we present NURBGen, the first framework for generating 3D CAD models from natural language using structured, symbolic NURBS representations. Unlike prior work that learns dense latent codes (Fan et al. 2024), we treat each NURBS surface as a language-aligned object: a sequence of tokens encoding control points, degrees, weights, and knot vectors in JSON format. This allows us to formulate text-to-CAD as a language modeling task. We fine-tune a large language model (Qwen3-4B) to map textual descriptions to these NURBS parameters, producing outputs that are editable and directly compatible to BRep format. To support this, we construct partABC, a curated dataset of more than 300k part-level CAD models from the ABC dataset, each represented as a sequence of NURBS surfaces and serialized with manageable context lengths ($\leq$ 8k tokens). We also generate high-quality natural language descriptions of the CAD models using an automatic annotation pipeline for the supervised fine-tuning task.
A key design choice to manage context length is our use of untrimmed NURBS surfaces similar to NeuroNURBS (Fan et al. 2024). However, this introduces a limitation it cannot capture trimmed geometry precisely. To address this, we propose a hybrid symbolic representation that replaces NURBS with primitive analytic curves (e.g., circles, B-splines, arcs, and lines) to accurately model such faces. This maintains the structural format required for LLM fine-tuning and inference. Our experiments demonstrate that NURBGen can outperform the current state-of-the-art methods in high-fidelity text-to-CAD generation as shown in Figure 1. Our contributions can be summarized as follows
@@
CAD Generation: Earlier approaches to CAD generation primarily focused on low-level geometry tasks such as surface fitting (Sharma et al. 2020; Liu et al. 2023), point cloud classification (Qi et al. 2017b, a), BRep segmentation (Dupont et al. 2022; Lee et al. 2023; Mallis et al. 2023), or BRep structure prediction (Guo et al. 2022; Ali et al. 2024), rather than generating fully parametric CAD models. A major shift came with DeepCAD (Wu et al. 2021), which introduced a design-history-based representation where CAD models are expressed as sequences of 2D sketches and 3D operations (e.g., extrusions). This formulation enabled sequence-to-sequence modeling of CAD generation. Building on this, later works explored cross-modal CAD synthesis from point clouds (Khan et al. 2024a; Dupont et al. 2024; Rukhovich et al. 2024), images (Chen et al. 2024), natural language (Li et al. 2025; Lv and Bao 2025; Khan et al. 2024b; Li et al. 2024; Govindarajan et al. 2025; Wang et al. 2025), or combinations thereof (Kolodiazhnyi et al. 2025; Xu et al. 2025). While design-history representations are highly interpretable and editable, their reliance on proprietary CAD operation data presents a major bottleneck for large-scale public research. Public datasets like DeepCAD-170k (Wu et al. 2021), Fusion360-8k (Willis et al. 2021), and CADParser-50k (Zhou et al. 2023) are limited in size and complexity, often consisting of simple, synthetic parts (Govindarajan et al. 2025), which restricts generalization to real-world scenarios. Alternative approaches for CAD generation operate directly on BRep geometry (Lambourne et al. 2021) or leverage SDF supervision (Ren et al. 2022; Li et al. 2023; Yu et al. 2022). However, these methods don’t generalizes well. Recent work BrepGen (Xu et al. 2024b), for example, generates BRep topology including vertices, edges, and faces via a hierarchical latent diffusion model. In contrast, we represent BReps as sequences of structured NURBS surfaces, allowing us to frame text-to-CAD as a language generation task. This enables fine-tuning an LLM on partABC, which is more diverse and larger than those used in prior work.
@@
Nurbs Modeling: The adoption of analytic surfaces like NURBS in learning-based systems remained limited (Böhm et al. 1984; Mykhaskiv et al. 2018). NURBSDiff (Prasad et al. 2022) introduced differentiable NURBS fitting for geometry optimization and reconstruction. (Worchel and Alexa 2023) proposed differentiable rendering of NURBS surfaces for inverse graphics tasks. The most relevant prior work is NeuroNURBS (Fan et al. 2024), which encodes untrimmed NURBS surfaces using a non-autoregressive transformer autoencoder into latent vectors for supporting tasks like reconstruction or segmentation, but not text-conditioned generation. However, the exclusive use of untrimmed NURBS surfaces limits generalization, as not all CAD models can be accurately represented without trimming. To address this, we adopt a hybrid strategy: while untrimmed NURBS serve as our primary representation, we replace them with analytic primitives such as lines, arcs, and B-splines for faces where NURBS fitting fails.
LLM for 3D Generation: LLMs have been widely adopted across domains such as robotics (Zeng et al. 2023) and 3D scene understanding or grounding (Xu et al. 2024a; Hong et al. 2023). Their application to 3D generation is a relatively new but promising direction, as LLMs offer strong spatial priors and multimodal reasoning capabilities. A central challenge, however, lies in encoding 3D geometry into a sequential format compatible with language modeling. Recent work such as LLaMA-Mesh (Wang et al. 2024) finetunes LLaMA (Grattafiori and et al 2024) to generate mesh vertices and faces as plain text, demonstrating the potential of autoregressive text-based 3D synthesis. In the CAD domain, existing LLM-driven text-to-CAD methods primarily rely on design-history-based representations (Li et al. 2025; Rukhovich et al. 2024; Xu et al. 2025; Zhang et al. 2025). However, these methods are constrained by the scarcity and simplicity of the public datasets. In contrast, our method introduces a structured NURBS-based representation that enables symbolic, surface-level generation. This formulation aligns naturally with language modeling and allows us to leverage the large-scale and geometrically diverse partABC dataset for fine-tuning LLMs for the text-to-CAD task.
@@ 1. CAD Representation
A BRep solid models geometry as a collection of topologically connected faces, each defined by a bounded parametric surface. In modern CAD systems, these surfaces are most commonly represented using NURBS surfaces due to their ability to accurately model both analytic primitives (e.g., planes, cylinders, tori) and complex free-form geometry with high continuity and compactness. To reconstruct a BRep solid in a symbolic and editable form, it is essential to extract the full set of NURBS surface parameters for each face. Using pythonOCC  [^2] https://github.com/tpaviot/pythonocc-core, we propose a robust pipeline for converting BReps into parametric NURBS representation.
@@
We represent each face using either its extracted NURBS parameters or its analytic primitive definition, based on reconstruction quality. In practice, about 70% of faces are modeled using NURBS, while 30% fall back to analytic primitives. This hybrid representation stored in structured JSON offers a more expressive and compact alternative to the purely NURBS-based format used in NeuroNURBS (Fan et al. 2024). While analytic primitives can represent simple geometry, they lack the flexibility to capture free-form surfaces. NURBS, on the other hand, provide a unified framework that can model both standard analytic shapes and complex free-form surfaces within a single patch—often replacing multiple primitives such as segmented arcs or partial cylinders. By combining both representations, our hybrid approach improves robustness and reduces parameter count for simpler shapes, resulting in shorter and more token-efficient inputs for LLM fine-tuning.
@@ 2. Annotation Pipeline
Supervised fine-tuning of our text-to-CAD model requires paired textual descriptions, but the ABC dataset lacks captions. To address this, we design an automated annotation pipeline using a VLM to generate high-quality captions for CAD models at scale as shown in Figure. 2 (middle).
Multi-View Rendering: Each BRep is first converted into a textureless triangular mesh and rendered from six viewpoints at a resolution of $512\times 512$ using Blender  [^3] https://docs.blender.org/api/3.6/index.html. Four of the camera views follow the orientation strategy proposed in (Sinha and Khan et al. 2025), while the remaining two capture the top and bottom perspectives. To enhance geometric perception and visual clarity, we enable Blender’s Freestyle renderer to overlay clean silhouette and edge contours on each image.
Metadata Guidance for Caption Generation: High-quality captions for CAD models should go beyond simple naming and incorporate essential geometric features such as the number of through-holes, overall dimensions, surface area, and volume. Previous work like Text2CAD (Khan et al. 2024b) leverages minimal JSON-based design history to guide vision-language models (VLMs), while MARVEL (Sinha and Khan et al. 2025) uses structured metadata for fine-grained 3D annotation. Building on these ideas, we extract geometric metadata that is often inaccessible to VLMs—specifically, length, width, height, surface area, volume, and the number of topological holes (genus). We compute overall dimensions by fitting an axis-aligned bounding box to the CAD geometry using OpenCascade’s Bnd_Box. Volume and surface area are obtained using OpenCascade’s built-in mass property computation (brepgprop.VolumeProperties and SurfaceProperties).
To estimate the number of through-holes, we first generate a watertight mesh from BRep. We then compute the Euler characteristic, $\chi=V-E+F$, where $V,E,F$ are the mesh vertices, edges, and faces. Using the Euler–Poincaré formula for closed 2-manifolds, the genus is given by $g=0.5\times(2-\chi)$, which corresponds to the number of topological through-holes (Cibulka et al. ). This metadata is then injected into the annotation prompt, which guides the VLM to generate captions with precise measurements.
Caption Generation: We use InternVL3-13B (Zhu et al. 2025), a multi-view VLM, which takes six rendered views of the CAD model along with the metadata-augmented annotation prompt as input. It then processes multi-view images of the CAD model simultaneously to generate a coherent and geometry-aware caption. Rather than focusing solely on object category names, we prioritize shape-centric descriptions that capture structural characteristics (”a bent tube..”,a flat cylindrical bolt with six holes..). The inclusion of dimensional metadata and hole counts further grounds the captions in precise geometric details, resulting in more informative and reliable annotations as shown in Figure 5.
@@
where $l_{1}=0.35,l_{2}=0.3,l_{3}=0.25,l_{4}=0.1$ are selected using empirical experiment on 100 samples. token_count refers to the size of tokens after tokenizing the JSON using Qwen3 Tokenizer (Yang et al. 2025), through_holes counts the number of holes that pass through the entire part, and bbox_diag is the length of the diagonal of the part’s axis-aligned bounding box. Based on $\texttt{w}(B)$, we categorize parts into simple ($\leq 0.12$), moderate (0.120.120.230.23), and complex (>0.23>0.23) tiers. From 3M extracted parts, we retain 10% simple, 50% moderate, and 40% complex models, forming the final partABC dataset of $\sim$300k high-quality samples as shown in Figure 6.
@@
Implementation Details: We fine-tune Qwen3-4B model (Yang et al. 2025) using AdamW (Loshchilov and Hutter 2019) with a learning rate of $5{\times}10^{-5}$ and linear warm-up. LoRA (Hu et al. 2021) is applied with rank $64$ and $\alpha{=}128$. Training runs for 180180k steps with batch size $1$ on $4{\times}$H200 GPUs over 3 days. The context window is $8192$ during training and 1414k during inference, with temperature $0.3$. On RTX 3090, the model achieves a generation throughput of $\sim 800$ tokens per second. Figure 2 (Right column) shows the finetuning task.
Baselines: We compare against strong open-source baselines for text-to-CAD generation. While recent models like CAD-LLaMA (Li et al. 2025) and CADFusion (Wang et al. 2025) report promising results, their implementations are not publicly available. Moreover, to the best of our knowledge, there are no open-source models capable of generating NURBS-based CAD representations from text. Hence, we focus our comparison with open-source methods, including Text2CAD (Khan et al. 2024b), DeepCAD (Wu et al. 2021), and GPT-4o. We use official pretrained weights for Text2CAD and retrain DeepCAD for 100 epochs following the Text2CAD protocol. GPT-4o is evaluated using 2-shot prompting with example caption–JSON pairs.
@@
Caption Quality: Since VLMs are prone to hallucinations (Liu et al. 2024), we perform an evaluation to assess the accuracy of our automatically generated captions. We randomly sample 1,000 captions and provide them along with the corresponding six rendered views and associated CAD metadata to GPT-4o. GPT-4o is tasked with verifying the correctness of each caption. We observe an estimated accuracy of $\sim 85\%$ across the sampled set. This indicates that our captioning pipeline produces high-quality and semantically meaningful descriptions.
@@
Despite its strong performance, our approach has certain limitations. Figure 7 illustrates a few representative failure cases. For instance, in response to complex prompts (e.g., “Two-story house with gabled roof…”), NURBGen struggles to capture fine-grained architectural structure. In rare cases, we also observe geometric artifacts such as self-intersections or topological inconsistencies, as seen in the second example. Additionally, NURBGen has difficulty reconstructing prompts with engraving text (third example).